OpenStudy (anonymous):

Are these functions odd, even, or neither? 1.f(x) = -x^6 + 3x^3 + 2x -7 (i think even) 2.f(x) = -x^4 + 2x^4 (i think neither)

jimthompson5910 (jim_thompson5910):

f(x) = -x^6 + 3x^3 + 2x -7 notice how the exponents are both odd and even this mix means the function is neither

jimthompson5910 (jim_thompson5910):

if you want an odd function, all exponents must be odd if you want an even function, all exponents must be even

jimthompson5910 (jim_thompson5910):

this trick only works with polynomials

OpenStudy (anonymous):

aah, okay, thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (zzr0ck3r):

also you can use the definition \(EVEN \iff f(-x) = f(x)\) \(ODD \iff f(-x) = -f(x)\) \(NEITHER \iff NOT \ ODD \ AND \ NOT \ EVEN\)

OpenStudy (zzr0ck3r):

wait, @jim_thompson5910 x^3+1 is not odd

jimthompson5910 (jim_thompson5910):

1 = 1x^0

jimthompson5910 (jim_thompson5910):

so x^3 + 1 is the same as x^3 + 1x^0 we have an odd exponent (3) and an even exponent (0)

jimthompson5910 (jim_thompson5910):

I never said f(x) = x^3 + 1 was an odd function